Question 9 Not yet answered B1 200, B2 800, and P3 10, at what age would you earn the highest wage? Points out o 1 (Hint: Find the value of Age when dAge Suppose that wage is a function of age: Wage-A + β2Age-β3Agelf dwage0 Flag question Select one: O a. 20 Ob. 40 c. 60 d. 780 Question 10 Not yet answeredwhere the population standard deviation is known. What is the upper bound of the Points out of 1 P Flag question A test resulted in a sample mean of3.7 and a standard error of 95% confidence interval for a two-sided test of the population mean (the critical value is 1.96)? Select one: a. 1.98 b. 3.9 c. 7.62 d. 8.45
Answer #1

Q9. The differentiated equation is,

dwage/age = 0

0= b2-2b3*age

0=800-2*10*Age

20Age=800

Age = 800/20=400

Q10. Option c

The formula is,

= mean+Critical value*Standard error

= 3.7+1.96*2

=7.62